40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14420     Entry time: Tue Jan 29 16:12:21 2019
Author: Chub 
Type: Update 
Category:  
Subject:  

The foam in the cable tray wall passage had been falling on the floor in little bite-sized pieces, so I investigated and found a fiber cable that had be chewed/clawed through.  I didn't find any droppings anywhere in the 40m, but I decided to bait an un-set trap and see if we'd find activity around it. There has been none so far.  If there is still none tomorrow, I will move the trap and keep looking for signs of rodentia.  At the moment, the trap is in a box in front of the double doors at the north end of the control room.  Next it will be place in the IFO room, up in the cable tray. 

gautam: the fiber that was damaged was the one from the LSC rack FiBox to the control room FiBox. So no DAFI action for a bit...

ELOG V3.1.3-